OpenStudy (anonymous):

whats the distance between parallel lines when given y=-2x+1 and y=-2x+16

OpenStudy (ageta):

2x+y-1=0 2x+y-16=0 the distance between these two parallel lines, I-16-(-1)I/√(2^2+1^2) =15/√5 =5*3/√5 =3√5 (ans) theory: two parallel lines, ax+by+c1=0 and ax+by+c2=0 the distance between these two lines is, Ic1-c2I/√(a^2+b^2)

OpenStudy (ageta):

does that help?

OpenStudy (ageta):